OpenStudy (anonymous):

Raul received a score of 70 on a history test for which the class mean was 70 with a standard deviation of 4. He received a score of 79 on a biology test for which the class mean was 70 with standard deviation 7. On which test did he do better relative to the rest of the class?

OpenStudy (perl):

change to z score z = (x - mu ) / sigma for the first test we have z = ( 70 - 70)/4=0 second test we have z = (79-70)/7 =1.285 so he did better on the second test

OpenStudy (perl):

relative to the rest of the class q